### v1

# Are All Marine Species Created Equal? Performance Disparities in Underwater Object Detection

Abstract
Underwater object detection is critical for monitoring marine ecosystems but poses unique challenges, including degraded image quality, imbalanced class distribution, and distinct visual characteristics. Not every species is detected equally well, yet underlying causes remain unclear. We address two key research questions: 1) What factors beyond data quantity drive class-specific performance disparities? 2) How can we systematically improve detection of under-performing marine species? We manipulate the DUO dataset to separate the object detection task into localization and classification and investigate the under-performance of the scallop class. Localization analysis using YOLO11 and TIDE finds that foreground-background discrimination is the most problematic stage regardless of data quantity. Classification experiments reveal persistent precision gaps even with balanced data, indicating intrinsic feature-based challenges beyond data scarcity and inter-class dependencies. We recommend imbalanced distributions when prioritizing precision, and balanced distributions when prioritizing recall. Improving under-performing classes should focus on algorithmic advances, especially within localization modules. We publicly release our code and datasets.
1 Introduction
Oceans are crucial for the regulation and function of natural water and carbon cycles, so much so that all life on Earth relies on healthy marine ecosystems. However these ecosystems are under threat from anthropogenic impacts including pollution, over-fishing and climate change [25]. This increases the urgency and importance of monitoring these environments.
The increasing use of Autonomous Underwater Vehicles (AUVs) and Remotely Operated Vehicles (ROVs) in recent years has yielded large quantities of underwater imagery [20]. Object detection plays an important role in analyzing this data, by identifying and precisely locating specific marine targets within an image. This enables monitoring ecosystem health through estimates of species distribution, size of targets, population density, signs of diseases and the presence/absence of indicator species [8].
Figure 1: Example underwater image from the DUO dataset [17], showing the four target classes (top row): Holothurian, Starfish, Scallop and Echinus. Left middle: the original class distribution of DUO shown by the lighter colored bars on top, and the DUO dataset with down-sampled, class-balanced distribution in the solid bars below. Left bottom: localization performance (mAP@0.5) on single-class versions of DUO. The lightened top bars indicate the localization performance for each class under the original distribution, and the solid colored bar below demonstrates there is still a clear performance difference between the classes, even when there is an equal number of training instances.
However, object detection in underwater imagery is significantly more challenging than for terrestrial images. Not only is the image quality often degraded by turbidity, blur and low contrast [32], but there also are certain visual characteristics and behavioral habits associated with different marine species that complicate their reliable detection. Image datasets of marine species may exhibit class imbalance, varying sizes of individuals, intra-class variation in appearance, background camouflaging, clustering of individuals, and dense habitats. Naturally, this leads to a discrepancy in the object detection performance for different classes of marine species (Fig. 1).
The Detecting Underwater Objects (DUO) dataset [17] is a combination of six earlier datasets, and contains real seafloor images of scallops, sea urchins (Echinus), sea cucumbers (Holothurian) and starfish [17], with an example image seen in Fig. 1. Prior underwater object detection (UOD) works [33, 31, 6, 32, 10, 34, 15] use this dataset as a benchmark, and demonstrate a clear trend in the relative performance of the different classes (Tab. 1). In these works, the authors attribute the significant under-performance of the scallop class in particular to the scarcity of instances in the training data [17, 33, 10] or the performance difference is not addressed at all.
Author $\text{AP}_{ho}$} $\text{AP}_{ec}$} $\text{AP}_{sc}$} $\text{AP}_{st}$} mAP Comment
IoU threshold 0.5:0.95
Liu et al. [17] 58.6 69.1 41.3 65.3 58.6 “smallest number of instances”
Xu et al. [31] 59.5 68.5 41.6 64.9 58.6
Dai et al. [6] 68.2 75.2 56.3 76.7 69.1
IoU threshold 0.5
Yuan et al. [32] 81.0 84.1 57.6 85.2 77.0
Zhao et al. [33] 85.3 93.9 64.0 95.2 84.6 “scarcity of scallop instances”
Feng et al. [10] 91.1 93.3 73.3 95.3 88.4 “scarcity of scallop targets”
Zhao et al. [34] 79.6 93.0 65.8 93.1 82.9
Li et al. [15] 84.7 92.6 70.8 93.3 85.4
Table 1: Results reported in previous studies on the DUO dataset. We show the per-class APs (refer to Sec. 3), for ho: holothurian, ec: echinus, sc: scallop, st: starfish, as well as the mAP. Cell color indicates AP values more than 5% above the respective mAP in green, AP values within ±5% in yellow, and AP values more than 5% below in red. There is persistent under-performance for the scallop class across all studies. We summarize the authors’ justification for the class discrepancies, noting that the scallop under-performance is often addressed purely qualitatively or omitted entirely.
Building on these observations, this work does not propose a new architecture but instead provides a systematic analysis of the under-performance of specific species in underwater object detection, moving beyond the commonly cited explanation of data scarcity. We investigate two important factors: the dataset characteristics, specifically the class distribution and data quantity; and the influence of inherent visual characteristics of the target objects. While the former can be changed when collecting the data and choosing the samples, the latter needs to be addressed on the algorithmic level. We also decompose the object detection architecture into the core tasks of localization and classification, to examine which stages are more impacted by these different data challenges. Understanding the underlying reasons for the class-wise detection performance is crucial to design improved object detectors for underwater ecosystems.
In this work, we present the following contributions:
  1. 1.
    We systematically decompose the object detection process into localization and classification stages to investigate the causes of class-specific performance disparities, finding that intrinsic visual features play a key role.
  2. 2.
    We conduct a dedicated localization study, using the TIDE toolkit [1] to analyze detection errors across different species, and identify that foreground-background discrimination is the main bottleneck, particularly for visually camouflaged species.
  3. 3.
    We study the classification stage of the object detector, revealing persistent performance gaps under balanced conditions. Further, we identify inter-class dependencies and a distribution-dependent tradeoff between precision and recall.
  4. 4.
    We highlight the importance of class-aware performance evaluation and provide recommendations to improve underwater object detector performance, considering common use cases and operational requirements.
We make our evaluation and data processing code available to foster further research in this area [^1] Link will be added upon acceptance..
2 Related Works
Object detection combines the basic tasks of localization, i.e. defining an object’s position using a bounding box, and classification, i.e. categorizing an object into a specified class, to find and recognize multiple objects within an image [22]. Here, we focus on Underwater Object Detection, and seek to establish the key factors which influence species-specific detection performance. To this end, we first review the unique characteristics and challenges of the underwater domain in Sec. 2.1, before summarizing recent underwater object detection approaches in Sec. 2.2, and describing methods and metrics for error evaluation in Sec. 2.3.
2.1 Challenges of the Underwater Domain
Transferring generic object detection to marine environments poses unique challenges caused by the characteristics of underwater imagery; these include poor visibility, turbidity, blur, low contrast, light scattering, and color distortion, and complicate feature extraction and foreground-background distinction [9, 25, 32, 8]. These challenges are compounded by the small size of many marine targets, frequent clustering of targets, lack of visual features, occlusion, and unclear object boundaries [8, 10, 15]. Few researchers consider the natural camouflage ability of marine species [8].
Learning the distinct features of marine species is therefore difficult and further complicated by in-class variation in appearance and scale, and differences in the angle, distance and speed of the camera [32, 25, 2]. While abundant underwater images can be obtained, the annotation effort is very time and cost-intensive and requires domain experts to reduce the risk of noisy and incorrect annotations [25, 2]. Finally, class imbalance is a common issue in underwater data, and is caused by differences in the economic value and rarity of certain species [18, 31].
This paper aims to separate the impacts of dataset characteristics i.e. quantity and class distribution, from the effects of species-specific visual features and behavioral traits.
2.2 Underwater Object Detection
Object detectors are typically transferred to the underwater domain as pre-trained detectors fine-tuned on marine datasets [2]. Popular approaches build upon Fast- and Faster-RCNN [11, 27] and YOLO [26]. Recent works have tailored detectors to the unique characteristics of the marine environment [2]. These include modified architectures with structural changes to the backbone for better feature extraction [32, 4, 34, 19, 10], improved neck structures for feature fusion, in particular focusing on small targets [34] and multiple scales [15], and adjusted loss functions, such as IoU-based [19, 10, 4] or intentional sample selection [3, 28]. Similarly, different training strategies were proposed to better understand complex underwater environments, prioritizing easy data in early training [3] or conversely, focusing on the most difficult samples with the highest loss values first [29].
To address the image quality issue, underwater image enhancement is a popular research direction [30]. However, several studies have already shown that it does not strictly correlate to better performance [6, 31, 33]. Other common data pre-processing includes data augmentation to diversify the training samples for better generalization [5]. Models can be specifically prepared for the underwater domain by simulating overlap, occlusion and blurring in training samples using RoIMix [16] or by changing position, size and quantity of objects to add minority class examples [18]. This directly addresses the common class imbalance problem in underwater data by synthetic oversampling.
While different methods allow majority classes to be under-sampled to create balance i.e. by randomly removing majority class samples in RandomUnderSampling (RUS[14], minority classes can be over-sampled by reusing or generating more examples. However, this increases the likelihood that duplicated samples provide insufficient information and cause overfitting [2]. Additionally, majority and minority class objects are often captured in the same image, so attempting to over-sample the minority class can also inadvertently increase majority class examples [5]. Alternatives to manual over/under-sampling include weighted sampling, in which underrepresented classes are assigned higher weights and importance during training [14].
2.3 Failure Analysis
The object detection performance on single classes is typically measured in Average Precision (AP), which is the area under the precision-recall curve (see Sec. 3), whereas the overall detector performance calculated by averaging the class APs to obtain the mean AP (mAP).
However, if only the aggregated metrics are used, the localization accuracy (i.e. how well bounding boxes are placed) and potential performance discrepancies across classes remain hidden. The respective influence of False Negative (FN) and False Positive (FP) errors are ignored. Hence different precision-recall curves can yield the exact same AP score, despite fundamentally different error patterns [23]. A detector might have low recall but high precision, making it suitable for risky tasks like autonomous underwater navigation. Conversely, another detector with the same mAP might instead provide high recall but low precision, which is preferred for ecological monitoring of invasive and harmful species, when it is most critical to avoid missing occurrences and enable early warnings.
Although these shortcomings are known and some researchers have developed alternative metrics [23, 7, 24], mAP is still the most commonly used metric, which is why we also report the mAP in this work.
Nonetheless, additional tools can be applied to gain further insights into the detector performance. Common tools include Diagnosis [13], TIDE [1] and False Negative Mechanisms [21]. While Diagnosis categorizes FPs and the FN Mechanisms focus on component failures causing FNs, TIDE proposes six main error types that can either result in FPs, in FNs or both, as explained and visualized in Fig. 2. To quantify the impact of each error type on overall performance, TIDE uses delta AP (dAP) by default, which measures the drop in AP caused by each error type in isolation. We utilize TIDE in our localization experiments to evaluate and report the common error types for each class.
Figure 2: Visualization of TIDE error types on an example DUO image. Ground truth is indicated in grey, predicted bounding boxes in color. The error types are: “Missed GT” (completely undetected ground truth, causes FN), “Classification - Cls” (correct box location but wrong class, causes FP & FN), “Classification and Localization” (wrong class and insufficient box alignment, causes FP & FN), “Background - Bkg” (background predicted as object, causes FP), “Duplicate - Dupe” (duplicate prediction for already matched ground truth, causes FP), and “Localization - Loc” (correct class but insufficient IoU overlap with ground truth, causes FP & FN). Classification, localization, and combined errors are formally counted as false positives in TIDE but also represent false negatives since they leave ground truths unmatched.
Despite the availability of failure analysis tools and known challenges of underwater object detection, existing studies primarily attribute class-specific performance disparities to data quantity alone [10, 17, 33], without systematic investigation of the underlying mechanisms. To our knowledge, no prior work has decomposed underwater object detection performance to isolate the relative contributions of classification versus localization errors across different marine species. In this work, we perform a systematic investigation of the key object detection stages, and perform failure analysis to isolate the key contributing factors of class-specific performance. This enables us to propose a comprehensive set of recommendations for underwater object detection in different operational scenarios.
3 Experimental Framework
This section details the general baseline for all our experiments, summarizing our approach, introducing the dataset of interest and the performance metrics used.
Approach: Object detection can be broadly considered as 3 key stages: (i) distinguishing between the background of an image and objects of interest; (ii) once detected as object of interest, tightly localizing it with a bounding box; and (iii) once localized as object of interest, classifying it between target classes. We investigate each of these in turn to identify how class-wise performance is affected by inherent visual features versus characteristics of the dataset. Steps (i) and (ii) are part of our localization analysis (Sec. 4), whereas (iii) is examined in a classification-focused study (Sec. 5). In every step we consider the influence of data quantity and distribution versus inherent class-specific properties.
Base Dataset: We perform our experiments on data from the DUO dataset [17]. DUO provides a standardized benchmark and assembles images from the Underwater Robot Professional Contest from 2017 to 2020, as well as the Underwater open-sea farm object Detection Dataset (UDD) [18]. In its original form, DUO provides 6,671 training and 1,111 test images, with 50,156 labeled echinus, 7,887 holothurian, 1,924 scallop and 14,548 starfish detections. We generate several subsets from the original DUO for our localization and classification studies, as detailed in Secs. 4.2 and 5.2.
Evaluation Metrics: We use mAP for the localization experiments and precision and recall for classification.
We use the common object detection metric AP for the localization task, with $\text{AP}=\int_{0}^{1}p(r)\,dr$. When averaged over the number of classes $C$, this becomes $\text{mAP}=\frac{1}{C}\sum_{i=1}^{C}\text{AP}_{i}$.
Additionally, we specifically evaluate bounding box accuracy per class by computing the center and area error. The center offset $O=\left\|\mathbf{c}_{p}-\mathbf{c}_{g}\right\|_{2}$ measures the Euclidean distance between the centers of the predicted ($\mathbf{c}_{p}$) and the ground truth bounding box ($\mathbf{c}_{g}$). The area error $E$ measures the relative deviation between predicted and ground truth bounding box areas as a percentage.
Precision $p=\frac{TP}{(TP+FP)}$ is the proportion of correct predictions (true positives $TP$) out of all predictions made. Recall, also known as the True Positive Rate (TPR), $\text{TPR}=\frac{TP}{(TP+FN)}$ is the proportion of correctly predicted objects out of all true objects. The False Negative Rate (FNR) is the proportion of missed targets out of the total, $\text{FNR}=\frac{FN}{(TP+FN)}$. We also consider the False Discovery Rate, which is the ratio of false positives to total predictions, $\text{FDR}=\frac{FP}{(FP+TP)}$. Finally, the F1 Score balances precision and recall: $F1=2\cdot\frac{\text{Precision}\cdot\text{Recall}}{\text{Precision}+\text{Recall}}$.
4 Localization Study
This section describes our localization analysis by explaining the methodology (Sec. 4.1), providing an overview of the datasets (Sec. 4.2), and detailing the implementation (Sec. 4.3). We then provide our results that show how well each class can be distinguished from the background (Sec. 4.4) and how well the object can be located with a bounding box (Sec. 4.5).
4.1 Method
In our localization experiments, we eliminate the classification stage interference by creating single-class versions of the DUO dataset. We train object detectors on individual classes to analyze class-specific localization challenges independently. The focus lies on assessing foreground-background confusion and examining bounding box quality. We also leverage TIDE [1] to gain detailed insights into the underlying error types.
We evaluate both the effect of data quantity and the impact of including or excluding unannotated images. We carefully consider possible operational requirements of underwater monitoring settings, and use these to inform the recommendations made.
4.2 Datasets
We obtain single-class DUO versions using the original DUO images but modify the annotations to omit the labels for all but one species at a time. A random 20% from the training split is selected for the validation split. To obtain further subsets, we select all “empty” images first and then add images containing the target species until reaching the desired instance count. Since images may contain multiple objects of the same class, we include all instances of the target species when an image is selected. The subsets include: (i) balanced sets in which we use the minority class (scallop) count as the target: 1,363 instances for training and 344 for validation, so that each single-class dataset contains an equal number of training objects; (ii) reduced sets with the instance count decreased to 75%, 50% and 25% of the balanced amount in each single-class dataset; and (iii) a special single-class DUO dataset (limited scallop) which consists only of images that actually contain the target of interest, scallop, for comparison purposes.
4.3 Implementation
Figure 3: Per class localization performance compared across datasets. From the left: mAP@0.5 for the single class DUO versions and balanced sets, which are then gradually reduced. Scallop performance is the same in the original DUO and balanced sets (all classes in the balanced set are downsampled to the scallop instance count), but still lower than all other classes, despite the same number of instances. Further reductions of the balanced set affect classes differently: starfish and echinus are impacted less by the decrease in samples, whereas holothurian and scallop classes exhibit significant reduction in the mAP.
For the localization experiments, we used Ultralytics YOLO11222https://github.com/ultralytics/ultralytics as object detector. The model was loaded with the default parameters, that is, 30 epochs, image resized to 640x640 pixels and data augmentations during training, including hue, saturation, brightness, translation, image scaling, horizontal flip and mosaic augmentation. Batch size was set to 8. The best model is saved based on the validation performance and evaluated on the test set with confidence threshold 0.25 and IoU set to 0.7. All localization experiments were run on an NVIDIA H100 GPU.
4.4 Results: Foreground-Background Separation
Our localization experiments reveal that class performance disparities persist even when the class distribution in the DUO dataset is balanced. As the left side of Fig. 3 illustrates, the single-class DUO versions show the drastic performance gaps that have motivated our work – mAP@0.5 scores for echinus and starfish greater than 90%, followed by holothurian at 82.5%, but substantially lower scallop performance at 44% – yet, even when all datasets are balanced to contain the same number of training labels, these performance gaps remain evident. The difference between scallop and holothurian for mAP@0.5 (the smallest gap) remains 27%, clearly demonstrating that the model struggles with scallop localization regardless of the quantity of training samples. This is the most major performance difference of all three detector tasks, and hints at challenges due to the inherent visual features.
Fig. 3 (right) shows that when the data quantity is gradually decreased in our balanced setup, the holothurian and scallop classes are more affected by the number of sample points in the training dataset and sensitive to data reductions. While the downwards trend generally highlights the benefits of larger training datasets, the varying rates of performance change again hint at inherent data characteristic differences.
single-class DUO balanced single-class sets
TPR FDR FNR TPR FDR FNR
Holothurian 76.74 21.89 23.26 56.81 19.24 43.19
Echinus 87.97 14.16 12.03 51.80 4.04 48.20
Starfish 88.02 16.95 11.98 68.32 8.42 31.68
Scallopall images 23.96 35.8 76.04 same as DUO
Scalloplimited 74.19 75.57 25.81
Table 2: True Positive Rate (TPR) or Recall, False Discovery Rate (FDR), and False Negative Rate (FNR) for each class under two training setups (refer to Sec. 3 for metric definitions). Cell colors indicate the relative contribution. Most of the scallop instances are undetected (76% FNR). Despite the increases in FN for Holothurian, Echinus and Starfish on the balanced set, their recall remains above the scallop value (24%). Training without background-only images (Scalloplimited row) increases the recall by 50.2%, but at the expense of an increase of 39.8% in the FDR.
(a) Holothurian DUO
(b) Echinus DUO
(c) Starfish DUO
(d) Scallop DUO = Balanced
(e) Holothurian Balanced
(f) Echinus Balanced
(g) Starfish Balanced
(h) Scallop Limited
Figure 4: Distribution of TIDE error types across all classes for both original and balanced DUO datasets. The imbalanced versions exhibit higher rates of FPs, and specifically background error, for Starfish, echinus and holothurian classes. In contrast, balanced datasets make the error profiles for these classes more similar to the scallop class, i.e. there are more missed ground truth detections. Only training on object-containing images (as seen in the (h) Scallop Limited sub-plot) reduces misses but massively increases background error.
Tab. 2 breaks down the true positive (recall), false discovery and false negative rates under imbalanced and balanced conditions. The results reveal that scallops generally suffer from extremely poor recall, with only 24% of true scallops being detected, compared to other classes with recall above 75% for single-class DUO. The balanced single-class sets resulted in decreased recall by 19.9%, 36.2% and 19.7% for holothurian, echinus and starfish, respectively. However, despite the equal number of training instances across the classes, scallops still show significantly lower recall at 24% than the other classes (51.8-68.3%), and are affected significantly more by FNs (76% for scallop compared with 31.7-48.2%), i.e. they are most often mistaken as part of the background. In the single-class DUO sets, the FDR is also highest for the scallop class at 35.8%, as compared to 14-21% for the other classes. When the sets are balanced, the FDR reduces by 2.7% to 19.2% for holothurian, by 10.1% to 4.0% for echinus, and by 8.5% to 8.4% for starfish, increasing the gap to the high scallop FDR to a maximum difference of 16.6%.
Qualitative inspection confirms the high background confusion, with scallops being the hardest to spot due to their ability to camouflage. In Fig. 2, the missed example is a small scallop blending into the background, resembling a faint light spot that could easily be mistaken for scattered light or a rock. The background error shows a rock falsely detected as a scallop, highlighting their visual similarity. In contrast, other classes are more distinct –echinus are dark and spiny, starfish have clear five-arm symmetry, and holothurians are elongated with surface spikes. Thus, scallops are arguably the most difficult class to distinguish from background clutter.
Overall, background confusion is the most prominent TIDE error type that occurs across all classes, as visible in Fig. 4. With the original DUO instances per class, starfish, echinus and holothurian still result in a considerable amount of background errors (causing FPs), the balanced versions result in high rates of missed GT errors. Of all the classes, scallops have the highest error rate relative to the number of test objects.
An important consideration when creating our datasets was whether to include background-only images. The comparisons between full scallop datasets and limited scallop-only datasets in Tab. 2 and Fig. 4 reveal that including diverse background images (even those without scallops) improves model generalization and heavily reduces false positives, though recall is reduced compared to training only on scallop-containing images.
4.5 Results: Bounding Box Placement
Tab. 3 indicates that once an object is distinguished from the background, further localization performance differences are not significant.
mAP@0.5 mAP@0.5:0.95
Class Value Deviation Value Deviation
Holothurian $0.7115$ 0.11-0.11 $0.5042$ 0.15-0.15
Echinus $0.7478$ 0.07-0.07 $0.5967$ 0.05-0.05
Scallop $0.4415$ 0.38-0.38 $0.3499$ 0.30-0.30
Starfish $0.8213$ $0.00$ $0.6507$ $0.00$
Table 3: The per-class mAP results of the balanced dataset and the deviation from the best‐performing class. The scallop class under-performance is less pronounced with higher IoU thresholds.
Specifically, when requiring better bounding box accuracy via a higher IoU threshold, the performance discrepancy becomes less between classes. This implies that localization precision is similar across classes, and the main cause of difference at lower thresholds is due to detection difficulty, not box alignment. This is underlined in Fig. 4: the balanced datasets have rarely any localization errors, i.e. if the object is not missed, the predicted bounding box usually is accurate enough. While there are some minor differences in the precision of bounding boxes across classes (Fig. 5), with the smaller and visually less distinct scallop objects having slightly larger center offsets and higher area errors more frequently, all classes follow a very similar error distribution regardless of imbalanced or balanced setup.
(a) BBox Center Offset
(b) BBox Area Error
Figure 5: Accuracy of predicted bounding boxes in balanced setup. (a) Bounding box center offset, with minor deviations but overall similar distribution across classes and datasets. (b) The bounding box area error, which highlights larger fluctuations for the small scallop objects but a similar trend overall.
5 Classification Study
This section describes our analysis on classification stage of object detection. Sec. 5.1 describes the methodology, before we explain the datasets (Sec. 5.2) and implementation (Sec. 5.3). Sec. 5.4 then highlights what factors impact how detected targets are assigned to the correct target class.
Precision Recall F1 Score
Dataset Version $\mathrm{Ho}$ $\mathrm{Ec}$ $\mathrm{Sc}$ $\mathrm{St}$ $\mathrm{Ho}$ $\mathrm{Ec}$ $\mathrm{Sc}$ $\mathrm{St}$ $\mathrm{Ho}$ $\mathrm{Ec}$ $\mathrm{Sc}$ $\mathrm{St}$
Classification DUO Reference Set 0.952 0.991 0.943 0.982 0.970 0.994 0.916 0.964 0.961 0.993 0.929 0.973
Imbalanced Minimum 0.884 0.978 0.776 0.959 0.897 0.988 0.703 0.921 0.890 0.983 0.733 0.940
Balanced Full 0.842 0.990 0.664 0.955 0.952 0.969 0.962 0.929 0.893 0.979 0.786 0.942
Classification DUO Reference Set + Weighted Sampling 0.950 0.993 0.946 0.987 0.979 0.995 0.934 0.966 0.964 0.994 0.940 0.976
Table 4: Classification precision, recall and F1 score (refer to Sec. 3 for definitions) across different dataset distributions. Results highlight a precision-recall tradeoff: balanced datasets improve scallop recall, imbalanced datasets result in higher precision. Weighted sampling mitigates scallop under-performance best.
5.1 Method
In the classification experiments, we focus purely on how well the query targets are assigned to the correct class. To isolate this stage from localization effects, we extract single-class image crops that simulate the output of a bounding box regressor and feed them directly to a classifier. This setup enables controlled manipulation of the training samples to establish the impact of data quantity and class imbalance.
For each experiment, we fine-tune a pre-trained classifier three times with different random seeds to minimize influence of random variation. Across experiments, the model architecture, hyperparameters, and test set remain consistent. We report results as the average across the three runs.
5.2 Datasets
Given that DUO was designed for object detection, we modified the data for our classification experiments to obtain single-object images. First, we cropped square images of every object with at least 10px padding around the bounding box. We randomly selected 20% of the training data as the validation set and removed all cropped images where the original bounding box overlapped by more than 10% with any other annotated object, thereby avoiding crops that contain multiple species. We call this cleaned subset the classification DUO reference set, which maintains the original imbalanced class distribution (69% echinus, 20% starfish, 9% holothurian, 2% scallop).
This serves as the basis for further subsets: (i) balanced sets with randomly down-sampled classes so that all provide the same data quantity as scallops, referred to as Balanced Full; (ii) reduced sets in which a) the total data quantity from the classification DUO reference set is decreased for an imbalanced minimum set of the same size as the balanced set maintaining class distribution, referred to as Imbalanced Minimum and b) the total data quantity from the balanced set is gradually decreased keeping equal instance count per class; and (iii) differently distributed sets where only samples of one class at a time are gradually reduced. We also conduct an experiment on the classification DUO reference set with weighted sampling to address the class imbalance.
5.3 Implementation
We used a ResNet-18 [12] pre-trained on ImageNet as our classifier using PyTorch. We fine-tune this model for 30 epochs with a batch size of 32. All images are resized to 224 x 224 pixels and normalized using ImageNet means and standard deviation. The training images are augmented with random horizontal flip and random rotation within $\pm$15 degrees. We use standard cross-entropy loss and Adam as optimizer with an initial learning rate of 0.001 that we reduce by the factor 0.5 every time no validation improvement has been recorded for 3 epochs, stopping at the minimum learning rate of 0.00003. All classification experiments were conducted on an Apple MacBook Pro M4 with MPS.
5.4 Results
Tab. 4 reveals a fundamental precision-recall tradeoff based on the training distribution. For scallops, imbalanced training yields higher precision but lower recall, while balanced training increases recall by 4.6% (from 91.6% to 96.2%) but significantly reduces precision by 27.9% from 94.3% to 66.4%. This pattern is most pronounced for scallops, with other classes showing greater robustness across training strategies. Scallop under-performance persists across all sets but is almost negligible when weighted sampling is applied. When the data is quantitatively balanced (Balanced Full), all classes achieve comparable and high recall but significant performance disparities are evident in precision: Fig. 6 reveals that even with balanced training, scallops achieve 55-70% precision versus >>95% for echinus and starfish, demonstrating intrinsic visual challenges.
This demonstrates that while there are benefits for larger training datasets, classification performance differences caused by inherent data characteristics will persist. The scallop class exhibits both the lowest overall precision, and the largest performance fluctuations as data is reduced, suggesting that scallop features are inherently less distinctive and more sensitive to data availability.
Figure 6: Per class classification performance on incrementally reduced balanced subsets (Solid lines: Precision; dashed lines: Recall). Recall is consistently high across all classes but there are significant gaps in precision, with the scallop class experiencing the largest fluctuations with reduced data.
Fig. 7 shows that reducing training data for other classes significantly impacts scallop classification performance, even when scallop training data remains constant. Interestingly, reducing starfish data to 30% causes an 8.2% drop in scallop precision and a 1.1% reduction in scallop recall, while reducing holothurian data results in a 7.9% precision decrease for scallops.
This interdependence suggests that the model learns scallop identity partially through negative examples, i.e. understanding what scallops are not. When other classes have insufficient representation, the model becomes less certain about class boundaries, leading to more false positive scallop predictions and reduced precision.
Figure 7: Effect of data reductions on performance of scallop class. From the classification reference set (100%), training samples are reduced to 30% one class at a time. Scallop precision declines as training data for other classes (particularly starfish and holothurian) is reduced, despite keeping scallop data constant. The model relies on sufficient negative examples to understand class boundaries.
6 Discussion and Conclusion
This work demonstrates that performance disparities in underwater object detection are not simply artifacts of imbalanced training data, but reflect fundamental differences in the visual characteristics and detectability of marine species. Through systematic decomposition of object detection into localization (foreground versus background followed by bounding box placement) and classification stages, we reveal that scallops consistently under-perform due to intrinsic visual properties that make them difficult to distinguish from seafloor backgrounds and other species.
Our key findings include: (1) Performance gaps persist under balanced training conditions, indicating intrinsic rather than data-driven challenges; (2) Class interdependence effects show that models learn target identity partly through negative examples from other species; (3) Training distribution creates fundamental precision-recall tradeoffs that should be aligned with operational requirements; and (4) Background discrimination is the primary bottleneck in underwater detection tasks.
These insights lead to practical recommendations: applications prioritizing comprehensive detection with high recall, such as conservation monitoring, should use balanced training. Commercial operations with high precision requirements should prefer imbalanced training. As separating targets from the background represents the primary technical challenge, research efforts should prioritize improving localization over classification accuracy via targeted algorithmic improvements or the use of extra sensing modalities.
By providing this systematic analysis and releasing our evaluation code and processed datasets, we aim to foster more targeted research approaches that address the specific challenges of underwater object detection rather than applying generic computer vision solutions.